# FRABench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v2

@@ 5.2 In-Domain Evaluation
Table 3: Average accuracy on FRA-OOD and FRA-OOD-H. “Seen UAs for Unseen Tasks” are used to assess task generalization, while “Unseen TAs for Unseen Tasks” evaluate aspect generalization.
5.3 Out-of-Domain Evaluation
5.3.1 Tasks Generalization Evaluation
We use FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to tasks. The results are shown in the regions outside the black sectors in Figure 6. GenEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Seen UAs for Unseen Tasks” column of Table 3, GenEval achieves overall average accuracies of 81% and 77.9% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. As illustrated in Figure 5a for aspect-level comparison, GenEval surpasses the baselines across most aspects, with only five aspects displayed here. We also provide several good cases in the Appendix G.
However, as shown in Table 3, results for IG are consistent across in-domain evaluation, with all baselines performing poorly. As shown in Figure 5b, all evaluators perform poorly on fidelity. Our analysis indicates that these aspects in IG require recognizing fine-grained object features, such as well-formed facial features in humans, which models fail to detect. GPT-4o exhibits similar limitations (detailed samples are shown in Appendix G). This suggests that, although the evaluator learns a broad range of aspects, it still lacks a better understanding of objective visual semantics.
5.3.2 Aspects Generalization Evaluation
We use the unseen TAs in FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L [31] tests on unseen aspects in Appendix B.2. The results are shown in the black sectors of Figure 6, and in Table 3. In terms of overall coverage, GenEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. It continues to demonstrate strong overall performance on the NLG, IU, and ITIG, with high consistency with both GPT-4o and human annotations, achieving overall average accuracies of 80.3% and 78.9%, respectively. For aspect-level comparison, as shown in Figure 5a, GenEval outperforms GPT-4o on unseen aspects such as suitability and explainability under NLG and IU in FRA-OOD-H. Additionally, for UAs generalization evaluation, since GenEval has already transferred relevant UAs from other tasks, it performs well on FRAUAs-OOD. Experimental results are provided in Appendix D.1. These results demonstrate that even without exposure to these unseen aspects during training, GenEval remains effective in assessment. This is largely due to the inherently interrelated nature of aspects, which enables the transfer of similar semantics and evaluation standards across tasks and modalities.
5.3.3 Public Benchmark Evaluation
Additionally, to verify the reliability and generalization of GenEval, we evaluate it on some public meta-evaluation benchmarks. Tables 4, 6, and6 show results for the NLG, IU, and IG tasks, respectively. GenEval outperforms Themis (NLG) and LLaVA-Critic (IU), both task-specific evaluators, indicating that our cross-task evaluator with aspect-level evaluation performs better than single-task evaluators. For IG tasks, GenEval performs on par with VisionReward and surpasses ImageReward. Overall, the results show that training on cross-modality and cross-task can lead to stronger performance than training on a single task or modality.
Method SummEval MANS
Coherence Consistency Fluency Relevance Ave. diff($\uparrow$)
@@ 5.3.3 Public Benchmark Evaluation
GPT-4o 58.0 64.2 79.1 85.1 64.3 72.8 60.1 67.1 65.3 72.3 68.5
Claude-3.5 63.5 70.6 81.6 87.9 73.7 83.1 59.6 66.6 69.6 77.0 68.4
Qwen2VL-72B 66.8 73.2 66.8 66.4 71.8 80.4 62.2 69.3 66.9 72.3 19.3
Qwen2VL-7B 48.8 52.5 35.5 30.6 44.8 49.8 41.7 44.5 42.7 44.3 60.2
Themis 60.7 62.1 81.8 86.0 73.3 77.7 54.4 54.6 67.5 70.1 44.2
Ours 64.6 71.8 75.2 83.5 74.7 84.5 61.3 67.6 69.0 76.9 69.3
Table 4: The results of GenEval compared with baselines on two benchmarks. “tau” indicates that accuracy is calculated with ties, and “diff” excludes tied pairs when calculating accuracy.
Method WildVision
tau ($\uparrow$) diff ($\uparrow$) $\tau$ ($\uparrow$)
GPT-4o 55.3 70.1 73.3
Claude-3.5 53.3 67.3 61.2
Qwen2VL-72B 67.3 59.6 65.5
Qwen2VL-7B 39.2 40.6 23.1
LLaVA-Critic 53.0 66.0 59.6
Ours 53.9 68.6 66.5
Table 5: For comparison in wild settings. $\tau$ indicates the comparison of model rankings using Elo ratings.
Method GenAI-Bench Winoground
tau($\uparrow$) diff($\uparrow$) Relation Object Both Ave.
@@
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4
Ours 53.6 65.5 57.5 59.1 80.7 65.7
Table 6: Comparison of GenEval and other baselines on two text-to-image meta-evaluation benchmarks.
5.4 Evaluation as Evaluator and Reward Model
In this section, we assess GenEval’s performance as an evaluator on the MT-Bench and MLLM-as-a-Judge benchmarks. We also assess its potential as a multimodal reward model using VLRewardBench. As shown in Table 8 and Table 8, GenEval outperforms Themis—the leading evaluator in the LLM-as-a-Judge setting. For MLLM-as-a-Judge, GenEval also demonstrates modest advantages over LLaVA-Critic and surpasses it in VLRewardBench. These results indicate that GenEval serves effectively as an evaluator for both NLG and IU tasks and further verify the potential of FRABench as training data for multimodal reward models.
Method MT-Bench
tau($\uparrow$) diff($\uparrow$)
GPT-4o 70.9 83.5
Claude-3.5 76.3 90.7
Qwen2VL-72B 75.9 88.7
Qwen2VL-7B 44.5 50.1
Themis 43.6 37.7
Ours 74.9 88.3
Table 7: Test results of the baselines as evaluators.
Method MLLM-as-a-Judge VLRewardBench
tau ($\uparrow$) diff ($\uparrow$) General Hallucination Reasoning Ave.
@@ 5.4 Evaluation as Evaluator and Reward Model
GPT-4o 58.1 67.0 50.2 81.4 74.8 68.8
Claude-3.5 58.4 68.3 38.5 82.6 66.1 62.4
Qwen2VL-72B 54.6 58.5 50.8 75.4 70.7 65.6
Qwen2VL-7B 41.3 44.6 45.1 62.8 62.5 56.8
LLaVA-Critic 55.6 65.5 42.0 41.2 60.0 47.7
Ours 57.2 67.0 46.4 57.7 71.1 58.4
Table 8: Test results of the baselines as multimodal evaluators and reward models on benchmarks.
6 Related Work
6.1 Coarse-grained Overall-oriented Evaluation
Previously, coarse-grained evaluators, which produce overall judgments based on one or several aspects, have been widely explored across various evaluation tasks [32, 33, 34, 35, 36, 37]. For instance, in NLG, Wang et al. [38] first introduce PandaLM, a fine-tuned LLM designed to evaluate pairwise texts based on several aspect. Similarly, Li et al. [39] propose Auto-J, an evaluator capable of handling a broader range of tasks and aspects, supporting both pointwise and pairwise evaluation settings. In image generation evaluation, Xu et al. [10] develop ImageReward, an evaluator trained on a large-scale human annotation dataset. Using alignment and fidelity as reference dimensions, ImageReward provides an overall assessment. Despite showing good overall performance, coarse-grained evaluations lack the granularity needed to diagnose specific model deficiencies and often introduce aspect bias into the evaluation process.
6.2 Fine-grained Aspect-level Evaluation
To address the limitations of coarse-grained evaluation, recent studies have shifted toward fine-grained evaluation [34, 40, 6, 13, 41, 42, 43], where LMs are fine-tuned on multi-aspect datasets to produce aspect-specific judgments. For example, in NLG, Hu et al.[9] propose Themis, an LLM trained on the GPT-4 annotated NLG-Eval corpus. In image understanding, LLaVA-Critic[12] is the first fine-grained evaluator integrating diverse criteria, showing strong correlation with GPT-4o. For image generation, Xu et al.[11] introduce VisionReward, a VQA-based evaluator that assesses image quality across fine-grained aspects. While these methods improve over coarse-grained evaluation by targeting specific aspects, they struggle with scalability across different tasks and aspects. Liu et al.[5] further investigate aspect generalization through X-Eval, a two-stage learning framework that incorporates auxiliary aspects and demonstrates partial generalization in NLG. However, its reliance on predefined reference aspects and lack of open-source implementation restricts its usability and reproducibility.
7 Conclusion, Limitation and Future Work
This paper introduces GenEval, a generalizable fine-grained evaluator for cross-task and cross-modal evaluation. We start by building a comprehensive aspect tree that leads to the creation of FRABench, a large-scale evaluation dataset. GenEval is trained on diverse tasks and aspects to enhance its generalization capabilities. Experimental results demonstrate that GenEval achieves strong generalization through aspect-level evaluation. However, compared to the other three tasks, GenEval and other LMMs have relatively limited performance in image generation. This may primarily be due to existing LMMs’ insufficient active visual semantic understanding, which fails to autonomously identify vision defects. In future work, we plan to explore RAG-based methods to address LMMs’ lack of visual grounding and further evaluate GenEval’s effectiveness in generating DPO training data and its potential with reinforcement learning using FRABench.
Appendix A Aspect Sources and Definitions
@@
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Table 13 to 26. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type, as defined in Figure 3. Additionally, underscores are used to benchmark extended TAs for ITIG.

### v3

@@ Introduction
Table 2: Average accuracy on FRA-OOD and FRA-OOD-H. Seen UAs for Unseen Tasks are used to assess task generalization, while Unseen TAs for Unseen Tasks evaluate aspect generalization.
Experimental Setup
Benchmarks.
(1) Out-of-Domain Evaluation: we use self-constructed FRA-OOD and FRA-OOD-H. (2) Evaluation as LLM-as-a-Judge: we select public meta-benchmarks from three types of tasks. For NLG, we use MT-Bench (Zheng et al. 2023), SummEval (Fabbri et al. 2021), and MANS (Guan et al. 2021). Since SummEval and MANS only provide scores, we generate pairwise samples for evaluation through sample pairing. For IU, we use WildVision (Lu et al. 2024), MLLM-as-a-Judge (Chen et al. 2024), and VLRewardBench (Li et al. 2024c). For IG, GenAI-Bench (Li et al. 2024a) Winoground (Thrush et al. 2022), and Pickapic (Kirstain et al. 2023) are selected. All meta-benchmarks contain human annotation. Moreover, we carefully check that the training set contains no overlapping samples with the meta-benchmarks to ensure fair comparison. Detailed descriptions of each meta-benchmark can be found in Appendix C. (3) Multi-Aspect and Multi-Task Assessment Learning: we use FRA-ID to investigate the multi-aspect learning and the meta-benchmarks to verify multi-task synergy.
Method SummEval MANS MT-Bench
@@
GPT-4o 58.0 64.2 79.1 85.1 64.3 72.8 60.1 67.1 65.3 72.3 68.5 70.9 83.5
Themis 60.7 62.1 81.8 86.0 73.3 77.7 54.4 54.6 67.5 70.1 44.2 43.6 37.7
Auto-J - - - - - - - - 60.4 67.0 68.2 73.0 85.5
Prometheus 2 55.2 62.1 65.5 74.7 61.6 69.8 55.0 61.8 59.3 67.1 69.0 55.1 72.0
Ours 64.6 71.8 75.2 83.5 74.7 84.5 61.3 67.6 69.0 76.9 69.3 74.9 88.3
Table 3: Evaluation as LLM-as-a-Judge for NLG task. “tau” indicates that accuracy is calculated with ties, and “diff” excludes tied pairs when calculating accuracy. Bold and underline indicate the first and second best results.
Method WildVision MLLM-as-a-Judge VLRewardBench
@@
GPT-4o 55.3 70.1 73.3 58.1 67.0 50.2 81.4 74.8 68.8
Claude-3.5 53.3 67.3 61.2 58.4 68.3 38.5 82.6 66.1 62.4
Qwen2VL-72B 50.3 59.6 65.5 54.6 58.5 50.8 75.4 70.7 65.6
LLaVA-Critic 53.0 66.0 59.6 55.6 65.5 42.0 41.2 60.0 47.7
Qwen2VL-7B 39.2 40.6 23.1 41.3 44.6 45.1 62.8 62.5 56.8
w/ IU. 47.3 60.1 58.3 50.1 59.0 45.6 57.6 68.5 57.2
w/ IU+ITIG. 51.2 65.2 64.1 54.1 64.8 46.5 56.9 70.7 58.0
w/ IU+IG. 52.7 67.9 66.0 56.1 66.8 46.4 57.0 71.0 58.1
Ours 53.9 68.6 66.5 57.2 67.0 46.4 57.7 71.1 58.4
Table 4: Evaluation as LLM-as-a-Judge for IU tasks. We evaluate baselines across three meta-benchmarks.
Baselines.
We experiment with representative models from two categories: (1) Prompting Models: GPT-4o, Claude-3.5, Qwen2-VL-72B-Instruct (Qwen2VL-72B), and Qwen2-VL-7B-Instruct (Qwen2VL-7B), implemented using the same prompts as UFEval. (2) Fine-tuned Evaluators: Themis-8B (Hu et al. 2024b) (Themis), Auto-J (Li et al. 2023b), Prometheus 2 (Kim et al. 2024a), LLaVA-Critic-7B (Xiong et al. 2024) (LLaVA-Critic), ImageReward (Xu et al. 2023a), and VisionReward (Xu et al. 2024). Notably, ImageReward and VisionReward use different evaluation paradigms from ours. Auto-J cannot assess specific aspects, and Prometheus 2 requires reference answers. Therefore, we use these methods only as baselines in the LLM-as-a-Judge evaluation.
Out-of-Domain Evaluation
Tasks Generalization Evaluation. For task generalization, we employ the seen UAs and unseen tasks from UFEval during testing to control for test variables. The results are shown in the regions outside the black sectors in Figure 3  [^2] We use aspect abbreviations here; the full names can be found in Appendix A.. Detailed scores for each aspect are reported in Appendix D. UFEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Tasks Generalization Evaluation” column of Table 2, UFEval achieves overall average accuracies of 85% and 83% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. We also provide several good and bad cases in Appendix G.
Aspects Generalization Evaluation. We use the unseen aspects of unseen tasks in FRA-OOD and FRA-OOD-H to evaluate UFEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L (Wang et al. 2022) tests on unseen aspects in Appendix B. The results are shown in the black sectors of Figure 3. In terms of overall coverage, UFEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. In the right column of Table 2, IUFEval continues to demonstrate strong performance on the NLG, IU, IG and ITIG, with high consistency with both GPT-4o and human annotators, achieving overall accuracies of 86.2% and 83.2%, respectively. These results demonstrate that even without exposure to unseen aspects during training, UFEval remains effective in assessment. This is largely due to the inherently interrelated nature of aspects, which enables the transfer of similar semantics and evaluation standards.
Evaluation as LLM-as-a-Judge
Method GenAI-Bench Winoground Pickapic
tau($\uparrow$) diff($\uparrow$) Relation Object Both Ave. tau($\uparrow$) diff($\uparrow$)
@@
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5 54.4 59.2
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1 49.1 53.7
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0 38.6 38.3
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4 48.9 58.0
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4 48.8 55.7
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6 38.1 40.2
w/ IG. 46.6 59.4 53.1 52.0 71.2 58.7 45.8 53.1
w/ IG+ITIG. 50.1 62.6 56.0 58.2 74.2 62.8 47.1 55.9
w/ IG+IU. 52.5 64.5 57.3 58.0 78.5 64.6 49.5 56.9
Ours 53.6 65.5 57.5 59.1 80.7 65.7 50.0 57.3
Table 5: Evaluation as LLM-as-a-Judge for IG task. We evaluate baselines across three meta-benchmarks.
Figure 4: Multi-aspect assessment learning results. We train models using different training set combinations and test them on three visual tasks in FRA-ID to validate the effectiveness of multi-aspect learning.
For NLG tasks. The experimental results, shown in Table 3, demonstrate the effectiveness of UFEval in NLG tasks. Specifically, in SummEval and MT-Bench, UFEval outperforms most baselines, with Claude being the sole exception. In the MANA benchmark, UFEval achieves the best performance, with an accuracy of 69.3%. These results highlight UFEval’s strong capability for text understanding evaluation compared to domain-specific models in this field.
For IU tasks. The results are presented in Table 4. We also evaluate model-level consistency on WildVision using Elo ratings and Kendall’s Tau ($\tau$) to compare model rankings (Xiong et al. 2024). Results demonstrate that UFEval outperforms LLaVA-Critic across all benchmarks. LLaVA-Critic focuses on respective datasets training, whereas UFEval uses joint learning across multiple tasks, which leads to relatively superior performance. More comprehensive MLLM-as-a-Judge results are provided in Appendix D.
For IG Evaluations. The experimental results are shown in Table 5. Our evaluator outperforms ImageReward in all benchmarks, achieving accuracies of 65.5%, 65.7%, and 57.3%, respectively. Even compared to the state-of-the-art image evaluator VisionReward, the performance gap is minimal—only 0.9% lower on GenAI-Bench and 0.3% lower on Winogrounded. These results demonstrate the promising capability of UFEval in the image generation task.
Multi-aspect and Multi-task Assessment Learning
The construction of UFEval is based on our intuitive argument: jointly learning to assess multiple visual aspects and tasks may foster a synergistic effect. Therefore, we explore the effectiveness of multi-aspect and multi-task learning on the evaluators, respectively. Specifically, we experiment with various training data configurations to train the model, analyzing the influence of multi-aspect and multi-task learning. For instance, for IU tasks, we design three training configurations to investigate the impact of multi-aspect and multi-task training: (1) learning solely on IU aspect-level assessment (w /IU), (2) jointly learning IU and ITIG aspect-level assessment (w /IU+ITIG), and (3) jointly learning IU and IG aspect-level assessment (w /IU+IG).
Multi-task Assessment Learning. The results are presented in Table 4 and 5. Our results demonstrate that multi-task learning outperforms single-task training in enhancing evaluator performance. For instance, as shown in Table 4, when evaluating IU tasks, models trained jointly on both IU and IG tasks achieve superior overall accuracy compared to those trained exclusively on IU. These findings underscore the significant advantages of exploiting shared representations and complementary knowledge across diverse visual tasks, ultimately yielding a more high-performing evaluator.
Multi-aspect Assessment Learning. The results presented in Figure 4 show that learning with multi-aspect from different tasks improves the UFEval’s performance across most aspects. For example, in the IG task, incorporating relevant aspects from the IU, such as alignment in caption generation and recognition accuracy, enhances the evaluation of aspects like object alignment in IG. Similar improvements are observed in IU and ITIG. These results highlight the benefits of leveraging multi-aspect data to enrich the evaluator’s understanding and shared aspect knowledge.
Conclusion
This paper proposes UFEval, the first unified fine-grained evaluator with task and aspect generalization. Specifically, we start by building a comprehensive aspect tree that leads to the creation of FRABench, a large-scale, multi-modal, and aspect-level evaluation dataset. FRABench spans four tasks, 28 sub-tasks, and 112 diverse aspects. We then fine-tune an LMM on FRABench to develop UFEval. Our experimental results demonstrate that joint learning across diverse visual tasks and aspects yields significant mutual benefits and generalization capabilities. In future work, we aim to leverage UFEval to automatically construct high-quality preference pair datasets from model outputs for DPO training.
Appendix A Appendix A: Aspect Sources and Definitions
@@
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Tables 10 to 23. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type. Additionally, underscores are used to benchmark extended TAs for ITIG. Abbreviations for all aspects are provided in the table.
Appendix B Appendix B: Analysis of the FRABench
In this section, we provide detailed information and comprehensive analysis about the FRABench.
More Detailed Information about the Dataset
We provide the information about sub-tasks for NLG and IU in Table 6, as this information is not shown in Figure 2 of the main text. For the tasks in the “Generated” column in Figure 2 of the main text, due to the absence of pairwise data, we use four LMMs with varying performance levels to generate pairwise responses for each query, as illustrated in Table 7.
Aspect Diversity Analysis
Following prior work (Wang et al. 2022; Honovich et al. 2022), we analyze the ROUGE-L distribution between aspects in the training set and the FRA-OOD, respectively. Specifically, we sample pairs of aspects from the training set and compute their ROUGE-L scores. The overall distribution is shown in Figure 9, which indicates that the selected aspects are distinct from one another, confirming the inclusion of diverse and novel aspects in the training set. Additionally, we measure the ROUGE-L scores between aspects by sampling one aspect from the training set and another from the FRA-OOD. As shown in Figure 10, aspects in the training set and those in the FRA-OOD do not overlap, demonstrating the validity of our out-of-domain evaluations.
Position Bias Analysis
We provide the counts of evaluation samples in the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, showing cases where Response 1 is better than, equal to, or worse than Response 2, as illustrated in Figures 6 to 8. As shown in Figure 6, the four tasks in the training set maintain a relatively consistent distribution. In Table 8, we also evaluate the position consistency of UFEval on the FRA-ID and FRA-OOD. We perform two inferences on the same sample, swapping the positions of the responses in the second inference, and calculate the average consistency, i.e., whether the results from the two inferences align. These experimental results demonstrate that UFEval achieves higher positional consistency compared to other baselines.
Figure 5: Statistics of pairwise data in the training set.
Figure 6: Statistics of pairwise data in the FRAUAs-OOD.
Figure 7: Statistics of pairwise data in the FRA-ID.
Figure 8: Statistics of pairwise data in the FRA-OOD.
Aspect Selection Across Different Sub-tasks
We provide detailed information on the aspects selected for the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, as presented in Table 9. Additionally, we include statistics regarding the number of evaluation labels used for training and testing in each sub-task. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o. The human-annotated data in IG is sourced from the ImageRewardDB dataset (Xu et al. 2023a).
Appendix C Appendix C: Experiment Detail
Fine-tuning Details of UFEval
In our experiments, we fine-tune Qwen2-VL-7B-Instruct on the training dataset using the Llama-Factory framework. We use 64 NVIDIA A100 GPUs, where each GPU processes one sample, and employ gradient accumulation with a factor of 2. The total batch size is 128. Additionally, we offload the optimizer parameters to the CPU to reduce GPU memory usage. The training process is carried out over 3 epochs. We set the learning rate to 1e-5, with 1% warmup steps. The learning rate follows a cosine decay schedule, gradually reducing as training progresses. UFEval is trained to predict pairwise rankings based on the criteria in the evaluation prompt, and provide detailed justifications for the assigned judgments. Standard cross-entropy loss is applied to both judgments and justifications.
Detailed Information of Public Meta-benchmarks
We provide detailed descriptions and experimental information for the public meta-benchmarks:
  • SummEval (Fabbri et al. 2021): SummEval evaluates LLMs on text summarization across four aspects: Fluency, Consistency, Coherence, and Relevance. The dataset contains 100 instructions per aspect, each with responses from 16 different LLMs and corresponding human ratings (totaling 6.4k annotated samples). For testing, we randomly generate 9 comparison pairs from the 16 responses per instruction, creating 3.6k evaluation samples.
  • MANS (Guan et al. 2021): MANS evaluates LLMs on the story generation task. Each instruction includes 5 different responses with human annotation, totaling 400 samples. For each instruction, we randomly select two responses to form pairwise data, resulting in 400 evaluation samples.
  • MT-Bench (Zheng et al. 2023): MT-Bench is specifically designed to evaluate LLMs’ capability as evaluators. The benchmark incorporates pairwise comparison data for 80 questions. Each pairwise data receives multiple human annotations, resulting in a total of 3k votes. To ensure high-quality evaluation samples, we select only those cases where all human annotators reach complete consensus, ultimately obtaining 684 evaluation samples.
  • WildVision (Lu et al. 2024): WildVision comprises 11k human-annotated preference relations among LMM response pairs. Each relation includes a question-image pair and two responses generated by different models, with a human-annotated preference (including ties). For testing, we randomly select 2k response pairs from the dataset using the same testing protocol as LLaVA-Critic.
  • MLLM-as-a-Judge (Chen et al. 2024): MLLM-as-a-Judge establishes a novel framework for assessing how closely model evaluations align with human judgments. The benchmark aggregates 17k multimodal evaluation instances (image-instruction-response triplets) spanning 14 diverse benchmarks and incorporating outputs from 6 different LMMs. Through systematic pairwise comparisons conducted by human evaluators, the dataset provides 5,719 carefully annotated judgment cases for analysis.
  • VLRewardBench (Li et al. 2024c): VLRewardBench is a comprehensive benchmark spanning general multimodal queries, visual hallucination detection, and complex reasoning tasks. Through an AI-assisted annotation pipeline combining sample selection with human verification, it curates 1,250 high-quality examples specifically designed to probe model limitations.
  • GenAI-Bench (Li et al. 2024a): GenAI-Bench is a comprehensive benchmark with 1.6k compositional prompts to evaluate text-to-visual generation, surpassing the size and difficulty of existing benchmarks. Additionally, it provides over 15k human ratings for multiple aspects to support research on vision-language alignment metrics.
  • Winoground (Thrush et al. 2022): Winoground is a benchmark for evaluating vision-language models’ ability to perform visio-linguistic compositional reasoning—the capacity to understand how the meaning of an image changes when the same words are rearranged in a caption. It was hand-curated by expert annotators and labeled with fine-grained tags, containing 400 samples.
  • Pickapic (Kirstain et al. 2023): Pickapic is a large dataset of text-to-image prompts and real users’ preferences over generated images. Authors create a web app that enables text-to-image users to generate images and specify their preferences, which constructs 500 test samples.
Aspects Used in the public meta-benchmark
Since UFEval requires aspects for evaluation, we assign one to each benchmark. Specifically, we select suitable aspects from the TAs and UAs, and for unseen tasks, we define appropriate aspects based on task properties. Tables 24 to 27 show the aspects used by UFEval, with blue indicating unseen aspects. GPT-4o, Claude-3.5, Qwen2VL-7B, and Qwen2VL-72B also follow the same aspect assignments.